OpenStudy (anonymous):

Find the general solution to the differential equation y´ = sec^2 (πt).

OpenStudy (anonymous):

\[\frac{dy}{dt}=\sec^2\pi t\\ \int dy=\int\sec^2\pi t~dt\\ y=\frac{1}{\pi}\tan\pi t+C\]

OpenStudy (anonymous):

If you don't see the equivalence right away between the integral and the general solution, consider a substitution: \[u=\pi t~~\Rightarrow~~du=\pi~dt~~\iff~~\frac{1}{\pi}du=dt\] so then \[\int \sec^2\pi t~dt=\frac{1}{\pi}\int\sec^2u~du=\frac{1}{\pi}\tan u+C=\frac{1}{\pi}\tan\pi t+C\]

OpenStudy (anonymous):

Thank you that helped a lot!